Jio is preparing for his way to school. His house is 30 km away. What time will he be there at school if he leaves his house at 6:30 a.m. and travels at an average speed of 40 km/h?
1
Expert's answer
2016-03-09T08:36:42-0500
We can find the time that Jio needs to reach the school from the formula:
t=s/v=(30 km)/(40 km⁄h)=0.75 h=0.75∙60 min=45 min.
Then, Jio will be at school at 6:30 a.m.+45 min=7:15 a.m.
Answer:
Jio will be at school at 7:15 a.m.